# Lost in the FoG: Pitfalls of Models for Large-Scale Hydrogen Distributions

### v4

@@ 3.3 Incorporating redshift-space distortions
Here, $\hat{\boldsymbol{z}}$ is the line of sight and $v_{z}$ represents the line-of-sight velocities. Density conservation implies that the real- and redshift-space density fields are related with
@@
Here, $f=d\ln D(a)/d\ln a$, with $D(a)$ defined as the linear growth rate (Dodelson & Schmidt, 2020). $u_{z}$ represents the renormalized line-of-sight velocities $u_{z}(\boldsymbol{r})=-v_{z}(\boldsymbol{r})/(aHf)$. $\boldsymbol{x}=\boldsymbol{r}-\boldsymbol{r}^{\prime}$ is the separation between two points. delimited-⟨$\langle$..$\rangle$ … ⟩ represents the ensemble average. The remaining term is the difference between local line-of-sight velocities $\Delta u_{z}=u_{z}(\boldsymbol{r})-u_{z}(\boldsymbol{r}^{\prime})$. For more information on Equation 6, see Taruya et al. (2010).
While Equation 6 is challenging to evaluate exactly, it can be simplified at linear scales. In this regime, the density and velocity fields are directly related via continuity ($\delta_{L}+\nabla\cdot\boldsymbol{v}_{L}=0$), leading to the famous Kaiser (1987) result:
@@
We can obtain the Kaiser term for tracer auto and cross-power spectra by substituting the tracer bias (Equation 2) into the continuity equation $\delta_{m,L}+\nabla\cdot\boldsymbol{v}_{L}=0$, yielding $\delta_{t,L}/b_{t}+\nabla\cdot\boldsymbol{v}_{L}=0$. The linear tracer velocity field is assumed to be the same as all matter since both are subject to the same potentials (Fry, 1996; Tegmark et al., 2002; Desjacques et al., 2010). This results in the following expressions for the Kaiser effect of the auto and cross-power spectra, respectively:
@@
We strictly use constant bias values since Equations 9-10 are only valid when both the velocity and density fields are linear. Of the two fields, the velocity field becomes nonlinear at larger scales because of its higher sensitivity to tidal effects (Scoccimarro, 2004; Jennings et al., 2011; Reid & White, 2011; Chan et al., 2012; Howlett et al., 2017). Tracer biases, conversely, are relatively insensitive to the velocity field, typically remaining constant in the linear regime for the density field (Seljak & Warren, 2004; Jullo et al., 2012). Thus, we can assume the bias is constant when using Equations 9-10. We show the contribution of the Kaiser term to the monopole of the ${\rm H}$i-galaxy cross-power spectrum with the red line in Fig. 1, which comprises a vertical translation of the power spectrum on all scales.
On smaller scales, the random motion of matter within virialized structures stretches the apparent clustering within halos along the line-of-sight. This phenomenon is known as the fingers-of-God (FoG) effect (Jackson, 1972) and suppresses redshift-space clustering on small scales. The FoG effect is typically modeled with a damping term that can take a variety of different forms. We follow the analysis of Sarkar & Bharadwaj (2018), who tested some common damping profiles for ${\rm H}$i and found a small preference for
@@
where $\sigma_{i}$ is the pairwise velocity dispersion (Peacock & Dodds, 1994; Park et al., 1994; Sheth et al., 2001). We emphasize that Equation 11 is purely phenomenological – $\sigma_{i}$ is determined by fitting to the power spectrum. Furthermore, Equation 11 assumes that $\sigma_{i}$ is scale-independent, which is known to be false (Scoccimarro, 2004; Slosar et al., 2006; Loveday et al., 2018) although perhaps justified if $\sigma_{i}$ is weakly scale-dependent over the scales of interest. The impact of the FoG term can be seen in the yellow line of Fig. 1, suppressing the power spectrum at $k\sim 0.3$ h cMpc-1 with the effect strengthening toward smaller scales.
In many models of RSDs, the FoG and Kaiser effects are handled separately which implicitly assumes that they are decoupled. However, both effects originate from the same line-of-sight displacements and therefore cannot be truly independent of each other (Scoccimarro, 2004; Hikage & Yamamoto, 2015). Furthermore, the mapping between real and redshift space in Equation 6 couples the Kaiser and FoG effects (Taruya et al., 2010). $\Delta u_{z}$ subtracts out large-scale averages in the velocity field, rendering the associated exponential term sensitive to small-scale effects like the random motion of matter within virialized objects. We speculate on how this RSD treatment affects the performance of the models in Section 6.2.
@@ 3.4 General model
Equations 12-13 are broadly representative of the various techniques used to model tracer power spectrum. For example, HOD models typically use the same Kaiser and FoG terms but determine the matter power spectrum and bias via integrals over various occupation properties as a function of halo mass, which requires some additional assumptions (for review, see Cooray & Sheth, 2002). Traditionally, the pairwise velocity dispersion is determined with a fit to the power spectrum, although some works have derived $\sigma_{\mathrm{HI}}$ self-consistently within the halo framework (e.g., Zhang et al., 2020; Padmanabhan et al., 2023). Conversely, LPT approaches modify the Kaiser term to include higher-order components (Scoccimarro, 2004; Percival & White, 2009; Taruya et al., 2010; Vlah et al., 2015). We discuss the potential impact of these higher-order terms in Section 6.2.
In Fig. 1, we compare the final ${\rm H}$i-galaxy cross-power spectrum model (Equation 13, pink line) with all ingredients to the TNG300 values (dotted line). The TNG300 power spectra are calculated by placing the ${\rm H}$i and galaxy distributions into a $800^{3}$ grid smoothed with a cloud-in-cell assignment scheme. Our results are converged with the sampling of the grid (appendix B in Osinga et al., 2024). The redshift-space positions of ${\rm H}$i and galaxies are computed using their velocities along an arbitrarily-chosen line of sight. We then compute the power spectra with pylians (Villaescusa-Navarro, 2024). This procedure is used for all power spectra calculated from TNG300 for the remainder of this work.
@@ 3.5 Simplifying assumptions
1) Tracer biases and correlation coefficients are constant. Tracer biases and correlation coefficients are constant in the linear regime, but some studies have shown that nonlinear effects can emerge on surprisingly large scales (Umeh et al., 2016; Osinga et al., 2024). These nonlinearities are often neglected in ${\rm H}$i intensity maps since they are expected to be small compared to measurement uncertainties on large scales. Nonetheless, some works add perturbations to the bias to incorporate some nonlinear terms, like $b(k)\approx b_{0}+b_{1}k$ (e.g., McDonald & Roy, 2009; Saito et al., 2014; Springel et al., 2018). In our work, we test the two extremes to capture the full range of behavior: $b_{c}$, a constant bias, and $b(k)$, a completely scale-dependent bias.
@@ 4.1 Bias
Figure 2: Top: Bias with respect to matter for various baryonic tracers. Blue (blue), red (red), and all-galaxies (gray) are shown with ${\rm H}$i (brown) at $z=0$ (left), $z=0.5$ (center), and $z=1$ (right). Large-scale bias values provided in Table 1. Bottom: The percentage error caused by assuming a constant bias with gray dotted lines at ±$\pm 10\%$. The bias for all baryonic tracers at most redshifts differ from the assumed constant value by $\geq 10\%$ by $k\sim 0.2-0.3$ h cMpc-1. The exceptions are the $z=0$ blue galaxy bias and $z=1$ ${\rm H}$i bias, which arise from their low occupation of massive halos offsetting nonlinearities.
@@
Figure 3: The redshift evolution of the bias for each baryonic tracer, with linear fits (dashed lines) to the data points to help visualize the trends. The precise values presented in these fits should be taken with caution with only three data points. Dotted lines show the expected bias evolution of a population only subject to gravitational effects (Fry, 1996). Tracers associated with star-formation (red, blue, ${\rm H}$i) deviate from the passive evolution more than the all-galaxy bias, suggesting that quenching processes play a significant role in how they trace the matter distribution.
@@
The ${\rm H}$i and galaxy biases from TNG300 are shown in the top row of Fig. 2. The bias scale-dependencies are influenced largely by two factors: the halo bias and halo occupation (Chen et al., 2021; Wang et al., 2021b). As explained in Section 3, nearly all ${\rm H}$i and galaxies reside in halos and thus reflect the shape of the halo bias. The second factor, halo occupation, modulates this behavior based on how each tracer occupies halos as a function of halo mass. Larger halos are known to cluster more strongly than smaller ones, leading to greater biases and nonlinearities both for them and the tracers that they tend to host (Sheth & Tormen, 1999; Cooray & Sheth, 2002). For example, red galaxies are known to favor older and massive halos, resulting in a greater amplitude and a sharper slope in the bias at all redshifts (Kaiser, 1984; Gao et al., 2005; Zehavi et al., 2011). Conversely, blue galaxies and ${\rm H}$i inhabit smaller, more isolated halos (Papastergis et al., 2013), resulting in smaller biases and nonlinearities (Wang et al., 2007). The remaining tracer bias in Fig. 2, all-galaxies, falls between that of blue and red galaxies, as expected from the combination of the two color subsamples.
@@
The tendency of ${\rm H}$i and blue galaxies to occupy small halos opposes the scale-dependent contributions from the halo bias, which generates some of the unique bias shapes in Fig. 2. For example, the $z=1$ ${\rm H}$i bias and $z=0$ blue galaxy bias are constant to surprisingly small scales because the halo bias and occupation effects coincidentally offset. The occupation effect grows with time, even becoming dominant at 0.1 $\lesssim k\lesssim 1$ h cMpc-1 in the ${\rm H}$i bias at $z<1$ and creating the “dip” in the ${\rm H}$i bias at those scales. These ${\rm H}$i bias shapes agree with other theoretical (Pénin et al., 2018), computational (Wolz et al., 2016; Spinelli et al., 2020; Wang et al., 2021b), and observational ($z=0$, Anderson et al., 2018) works on ${\rm H}$i distributions. We note however that the amplitude of the ${\rm H}$i bias can vary substantially between the different works (further discussed in Section 6.1).
@@
Significant deviations from passive evolution suggest the influence of additional processes on the tracer’s clustering beyond gravity, such as changes in number density through mergers or satellite disruption (Bell et al., 2004; Guo et al., 2013; Skibba et al., 2014). We compare the redshift evolution of the various tracer biases to the expected passive evolution in Fig. 3. We fit a linear model to the three data points, with the slope as the only free parameter and assuming the $z=0$ value as the intercept.
@@
The key conclusion from Fig. 3 is that tracers sensitive to star-formation deviate more from passive evolution than tracers independent of star-formation. This trend arises primarily from changes in how star-formation dependent tracers occupy halos. At an earlier redshift, the star-forming galaxies that occupy the most massive, clustered halos also tend to be closest to quenching. Once these galaxies quench at a later redshift, the star-forming galaxies lose their most clustered component, inducing weaker clustering on average. However, these recently-quenched galaxies join the quiescent population as their least massive, clustered members, also suppressing clustering. Therefore, galaxy quenching reduces the clustering of all tracers dependent on star-formation, like ${\rm H}$i and blue and red galaxies (for more details, see Osinga et al., 2024). The stronger role of galaxy formation physics on the distribution of ${\rm H}$i and blue galaxies at low redshifts suggests that they may be poorer tracers of the matter distribution than a star-formation independent population.
@@ 4.2 Correlation coefficients
The small-scale decoupling is weakest for ${\rm H}$i and blue galaxies. Star-forming blue galaxies tend to be gas-rich and occupy the same regions of space as ${\rm H}$i, and thus strongly correlate even on small scales (Bigiel et al., 2008; Huang et al., 2012; Kauffmann et al., 2013; Papastergis et al., 2013; Hearin et al., 2016; Anderson et al., 2018). Conversely, red galaxies weakly correlate with ${\rm H}$i because they tend to inhabit gas-poor regions. ${\rm H}$i $\times$ Galaxy resides between ${\rm H}$i $\times$ Red and ${\rm H}$i $\times$ Blue, effectively representing their weighted average. A larger fraction of galaxies are quenched at later redshifts such that ${\rm H}$i $\times$ Galaxy approaches ${\rm H}$i $\times$ Red with time (Nelson et al., 2018). This quenching effect does not manifest in the large-scale values of ${\rm H}$i $\times$ Blue and ${\rm H}$i $\times$ Red, although they are reduced on small scales. Table 2 quantifies both the redshift and color trends with estimates for the constant correlation coefficients and $k_{10\%}$(definition in Section 4.1).

### v5

@@ 3.3 Incorporating redshift-space distortions
Here, $\hat{\boldsymbol{z}}$ is the line of sight and $v_{z}$ represents the line-of-sight velocities. Density conservation [^2] Strictly speaking, intensity mapping considers the specific intensity, not density, but the derivation follows a similar procedure regardless (S. Bharadwaj et al., 2001; S. Bharadwaj & S. S. Ali, 2004). implies that the real- and redshift-space density fields are related with
@@
Here, $f=d\ln D(a)/d\ln a$, with $D(a)$ defined as the linear growth rate (S. Dodelson & F. Schmidt, 2020). $u_{z}$ represents the renormalized line-of-sight velocities $u_{z}(\boldsymbol{r})=-v_{z}(\boldsymbol{r})/(aHf)$. $\boldsymbol{x}=\boldsymbol{r}-\boldsymbol{r}^{\prime}$ is the separation between two points. $\langle$$\rangle$ represents the ensemble average. The remaining term is the difference between local line-of-sight velocities $\Delta u_{z}=u_{z}(\boldsymbol{r})-u_{z}(\boldsymbol{r}^{\prime})$. For more information on Equation 6, see A. Taruya et al. (2010).
While Equation 6 is challenging to evaluate exactly, it can be simplified at linear scales. In this regime, the density and velocity fields are directly related via continuity ($\delta_{L}+\nabla\cdot\boldsymbol{v}_{L}=0$), leading to the famous N. Kaiser (1987) result:
@@
We can obtain the Kaiser term for tracer auto and cross-power spectra by substituting the tracer bias (Equation 2) into the continuity equation $\delta_{m,L}+\nabla\cdot\boldsymbol{v}_{L}=0$, yielding $\delta_{t,L}/b_{t}+\nabla\cdot\boldsymbol{v}_{L}=0$. The linear tracer velocity field is assumed to be the same as all matter since both are subject to the same potentials (J. N. Fry, 1996; M. Tegmark et al., 2002; V. Desjacques et al., 2010). This results in the following expressions for the Kaiser effect of the auto and cross-power spectra, respectively:
@@
We strictly use constant bias values since Equations 9-10 are only valid when both the velocity and density fields are linear. Of the two fields, the velocity field becomes nonlinear at larger scales because of its higher sensitivity to tidal effects (R. Scoccimarro, 2004; E. Jennings et al., 2011; B. A. Reid & M. White, 2011; K. C. Chan et al., 2012; C. Howlett et al., 2017). Tracer biases, conversely, are relatively insensitive to the velocity field, typically remaining constant in the linear regime for the density field (U. Seljak & M. S. Warren, 2004; E. Jullo et al., 2012). Thus, we can assume the bias is constant when using Equations 9-10. We show the contribution of the Kaiser term to the monopole of the $\rm H$}i-galaxy cross-power spectrum with the red line in Fig. 1, which comprises a vertical translation of the power spectrum on all scales.
On smaller scales, the random motion of matter within virialized structures stretches the apparent clustering within halos along the line-of-sight. This phenomenon is known as the fingers-of-God (FoG) effect (J. C. Jackson, 1972) and suppresses redshift-space clustering on small scales. The FoG effect is typically modeled with a damping term that can take a variety of different forms. We follow the analysis of D. Sarkar & S. Bharadwaj (2018), who tested some common damping profiles for $\rm H$}i and found a small preference for
@@
where $\sigma_{i}$ is the pairwise velocity dispersion (J. A. Peacock & S. J. Dodds, 1994; C. Park et al., 1994; R. K. Sheth et al., 2001). We emphasize that Equation 11 is purely phenomenological – $\sigma_{i}$ is determined by fitting to the power spectrum. Furthermore, Equation 11 assumes that $\sigma_{i}$ is scale-independent, which is known to be false (R. Scoccimarro, 2004; A. Slosar et al., 2006; J. Loveday et al., 2018) although perhaps justified if $\sigma_{i}$ is weakly scale-dependent over the scales of interest. The impact of the FoG term can be seen in the yellow line of Fig. 1, suppressing the power spectrum at $k\sim 0.3$ h cMpc-1 with the effect strengthening toward smaller scales.
In many models of RSDs, the FoG and Kaiser effects are handled separately which implicitly assumes that they are decoupled. However, both effects originate from the same line-of-sight displacements and therefore cannot be truly independent of each other (R. Scoccimarro, 2004; C. Hikage & K. Yamamoto, 2015). Furthermore, the mapping between real and redshift space in Equation 6 couples the Kaiser and FoG effects (A. Taruya et al., 2010). $\Delta u_{z}$ subtracts out large-scale averages in the velocity field, rendering the associated exponential term sensitive to small-scale effects like the random motion of matter within virialized objects. We speculate on how this RSD treatment affects the performance of the models in Section 6.2.
@@ 3.4 General model
Equations 12-13 are broadly representative of the various techniques used to model tracer power spectrum. For example, HOD models typically use the same Kaiser and FoG terms but determine the matter power spectrum and bias via integrals over various occupation properties as a function of halo mass, which requires some additional assumptions (for review, see A. Cooray & R. Sheth, 2002). Traditionally, the pairwise velocity dispersion is determined with a fit to the power spectrum, although some works have derived $\sigma_{\mathrm{HI}}$ self-consistently within the halo framework (e.g., J. Zhang et al., 2020; H. Padmanabhan et al., 2023). Conversely, LPT approaches modify the Kaiser term to include higher-order components (R. Scoccimarro, 2004; W. J. Percival & M. White, 2009; A. Taruya et al., 2010; Z. Vlah et al., 2015). We discuss the potential impact of these higher-order terms in Section 6.2.
In Fig. 1, we compare the final $\rm H$}i-galaxy cross-power spectrum model (Equation 13, pink line) with all ingredients to the TNG300 values (dotted line). The TNG300 power spectra are calculated by placing the $\rm H$}i and galaxy distributions into a $800^{3}$ grid smoothed with a cloud-in-cell assignment scheme. Our results are converged with the sampling of the grid (appendix B in C. K. Osinga et al., 2024). The redshift-space positions of $\rm H$}i and galaxies are computed using their velocities along an arbitrarily-chosen line of sight. We then compute the power spectra with pylians (F. Villaescusa-Navarro, 2024). This procedure is used for all power spectra calculated from TNG300 for the remainder of this work.
@@ 3.5 Simplifying assumptions
1) Tracer biases and correlation coefficients are constant. Tracer biases and correlation coefficients are constant in the linear regime, but some studies have shown that nonlinear effects can emerge on surprisingly large scales (O. Umeh et al., 2016; C. K. Osinga et al., 2024). These nonlinearities are often neglected in $\rm H$}i intensity maps since they are expected to be small compared to measurement uncertainties on large scales. Nonetheless, some works add perturbations to the bias to incorporate some nonlinear terms, like $b(k)\approx b_{0}+b_{1}k$ (e.g., P. McDonald & A. Roy, 2009; S. Saito et al., 2014; V. Springel et al., 2018). In our work, we test the two extremes to capture the full range of behavior: $b_{c}$, a constant bias, and $b(k)$, a completely scale-dependent bias.
@@ 4.1 Bias
Figure 2: Top: Bias with respect to matter for various baryonic tracers. Blue (blue), red (red), and all-galaxies (gray) are shown with $\rm H$}i (brown) at $z=0$ (left), $z=0.5$ (center), and $z=1$ (right). Large-scale bias values provided in Table 1. Bottom: The percentage error caused by assuming a constant bias with gray dotted lines at $\pm 10\%$. The bias for all baryonic tracers at most redshifts differ from the assumed constant value by $\geq 10\%$ by $k\sim 0.2-0.3$ h cMpc-1. The exceptions are the $z=0$ blue galaxy bias and $z=1$ $\rm H$}i bias, which arise from their low occupation of massive halos offsetting nonlinearities.
@@
Figure 3: The redshift evolution of the bias for each baryonic tracer, with linear fits (dashed lines) to the data points to help visualize the trends. The precise values presented in these fits should be taken with caution with only three data points. Dotted lines show the expected bias evolution of a population only subject to gravitational effects (J. N. Fry, 1996). Tracers associated with star-formation (red, blue, $\rm H$}i) deviate from the passive evolution more than the all-galaxy bias, suggesting that quenching processes play a significant role in how they trace the matter distribution.
@@
The $\rm H$}i and galaxy biases from TNG300 are shown in the top row of Fig. 2. The bias scale-dependencies are influenced largely by two factors: the halo bias and halo occupation (Z. Chen et al., 2021; Z. Wang et al., 2021). As explained in Section 3, nearly all $\rm H$}i and galaxies reside in halos and thus reflect the shape of the halo bias. The second factor, halo occupation, modulates this behavior based on how each tracer occupies halos as a function of halo mass. Larger halos are known to cluster more strongly than smaller ones, leading to greater biases and nonlinearities both for them and the tracers that they tend to host (R. K. Sheth & G. Tormen, 1999; A. Cooray & R. Sheth, 2002). For example, red galaxies are known to favor older and massive halos, resulting in a greater amplitude and a sharper slope in the bias at all redshifts (N. Kaiser, 1984; L. Gao et al., 2005; I. Zehavi et al., 2011). Conversely, blue galaxies and $\rm H$}i inhabit smaller, more isolated halos (E. Papastergis et al., 2013), resulting in smaller biases and nonlinearities (Y. Wang et al., 2007). The remaining tracer bias in Fig. 2, all-galaxies, falls between that of blue and red galaxies, as expected from the combination of the two color subsamples.
@@
The tendency of $\rm H$}i and blue galaxies to occupy small halos opposes the scale-dependent contributions from the halo bias, which generates some of the unique bias shapes in Fig. 2. For example, the $z=1$ $\rm H$}i bias and $z=0$ blue galaxy bias are constant to surprisingly small scales because the halo bias and occupation effects coincidentally offset. The occupation effect grows with time, even becoming dominant at 0.1 $\lesssim k\lesssim 1$ h cMpc-1 in the $\rm H$}i bias at $z<1$ and creating the “dip” in the $\rm H$}i bias at those scales. These $\rm H$}i bias shapes agree with other theoretical (A. Pénin et al., 2018), computational (L. Wolz et al., 2016; M. Spinelli et al., 2020; Z. Wang et al., 2021), and observational ($z=0$, C. J. Anderson et al., 2018) works on $\rm H$}i distributions. We note however that the amplitude of the $\rm H$}i bias can vary substantially between the different works (further discussed in Section 6.1).
@@
Significant deviations from passive evolution suggest the influence of additional processes on the tracer’s clustering beyond gravity, such as changes in number density through mergers or satellite disruption (E. F. Bell et al., 2004; H. Guo et al., 2013; R. A. Skibba et al., 2014). We compare the redshift evolution of the various tracer biases to the expected passive evolution in Fig. 3. We fit a linear model to the three data points, with the slope as the only free parameter and assuming the $z=0$ value as the intercept.
@@
The key conclusion from Fig. 3 is that tracers sensitive to star-formation deviate more from passive evolution than tracers independent of star-formation. This trend arises primarily from changes in how star-formation dependent tracers occupy halos. At an earlier redshift, the star-forming galaxies that occupy the most massive, clustered halos also tend to be closest to quenching. Once these galaxies quench at a later redshift, the star-forming galaxies lose their most clustered component, inducing weaker clustering on average. However, these recently-quenched galaxies join the quiescent population as their least massive, clustered members, also suppressing clustering. Therefore, galaxy quenching reduces the clustering of all tracers dependent on star-formation, like $\rm H$}i and blue and red galaxies (for more details, see C. K. Osinga et al., 2024). The stronger role of galaxy formation physics on the distribution of $\rm H$}i and blue galaxies at low redshifts suggests that they may be poorer tracers of the matter distribution than a star-formation independent population.
@@ 4.2 Correlation coefficients
The small-scale decoupling is weakest for $\rm H$}i and blue galaxies. Star-forming blue galaxies tend to be gas-rich and occupy the same regions of space as $\rm H$}i, and thus strongly correlate even on small scales (F. Bigiel et al., 2008; S. Huang et al., 2012; G. Kauffmann et al., 2013; E. Papastergis et al., 2013; A. P. Hearin et al., 2016; C. J. Anderson et al., 2018). Conversely, red galaxies weakly correlate with $\rm H$}i because they tend to inhabit gas-poor regions. $\rm H$}i $\times$ Galaxy resides between $\rm H$}i $\times$ Red and $\rm H$}i $\times$ Blue, effectively representing their weighted average. A larger fraction of galaxies are quenched at later redshifts such that $\rm H$}i $\times$ Galaxy approaches $\rm H$}i $\times$ Red with time (D. Nelson et al., 2018). This quenching effect does not manifest in the large-scale values of $\rm H$}i $\times$ Blue and $\rm H$}i $\times$ Red, although they are reduced on small scales. Table 2 quantifies both the redshift and color trends with estimates for the constant correlation coefficients and $k_{10\%}$(definition in Section 4.1).